OpenStudy (anonymous):

Evaluate the expression. r = <5, 9, -4>, v = <3, 3, -8>, w = <9, -6, -1> v - w

OpenStudy (mathmate):

if p=<a,b,c>, and q=<d,e,f> p-q=<a-d, b-e, c-f>

OpenStudy (anonymous):

I did that, and I don't if it's because I'm exhausted I got <27,-18,-8>

jimthompson5910 (jim_thompson5910):

@cariestrada you multiplied instead of subtracted

OpenStudy (anonymous):

ahhh jeez so it's <-6,9,-9> ?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

the issue is the last term -8 - (-1) = -9 is false

OpenStudy (anonymous):

-7!

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

But those are none of my answer choices :(

jimthompson5910 (jim_thompson5910):

what are your answer choices?

OpenStudy (anonymous):

<-15,-27,-21> 0 17 <-27, -18,8>

jimthompson5910 (jim_thompson5910):

could it be that they want the dot product instead of v - w ?

OpenStudy (anonymous):

That's what I was thinking, but the question is literally like I posted above, and it's a subtraction sign

jimthompson5910 (jim_thompson5910):

because when I did the dot product of v and w, I got v dot w = 3*9+3*(-6)+(-8)*(-1) = 27 - 18 + 8 = 17 which is a listed answer choice

jimthompson5910 (jim_thompson5910):

could be a typo on their end

OpenStudy (anonymous):

True, I'm gonna go with that. Thank you!

jimthompson5910 (jim_thompson5910):

no problem